# Bohr-Sommerfeld type surgeries and Disk Potential

### v1

@@ 6.1. Background on Neck-Stretching
Definition 6.2 (T-double-stretched manifold).
@@
where $\sim^{1}$ is the pair of identifications $\phi(\{a^{\prime}-\epsilon\}\times Z)=\{-T\}\times Z$ and $\phi(\{a^{\prime}+\epsilon\}\times Z)=\{T\}\times Z$ and $\sim^{2}$ is the similar pair of identifications at $\{b^{\prime}\pm\epsilon\}\times Z$. We define a T-double-stretched almost complex structure $J^{T}$ by extending the cylindrical almost complex structure naturally to $[-T,T]\times Z$.
@@
$\widehat{M}_{out-},\underbrace{M_{-l}:=\mathbb{R}\times Z,\cdots,M_{-1}\mathbb{R}\times Z}_{l\text{ times}},{M}_{in}:=\mathbb{R}\times Z,\underbrace{M_{1}:=\mathbb{R}\times Z,\cdots,M_{k}:=\mathbb{R}\times Z}_{k\text{ times}},\widehat{M}_{out+}$
where $\widehat{M}_{out}$ is obtained by attaching natural symplectization ends to $M\setminus\phi([a,b]\times Z).$ The manifolds ×$\mathbb{R}\times Z$, apart from $M_{in}$, are called the neck pieces. The collection of $l$ neck pieces are called the lower neck pieces and the collection of $m$ neck pieces are called upper neck pieces.
In case $\mathbb{Z}$ was a separating hypersurface, we call the two disjoint components of $\widehat{M}_{out}$ as $\widehat{M}_{out-}$ and $\widehat{M}_{out+}$ depending on which end of symplectization is added.
@@
If $u_{n}:\mathbb{D}^{2}\rightarrow(M^{n},J^{n},L^{n})$ is a sequence of $J^{n}$ holomorphic disc to the $n-$double-stretched manifolds, with boundary on $L^{n}$, and with bounded symplectic energy, i.e. $E(u)=<E<\infty$, then there is a subsequence of $u_{n}$ that converges to a stable broken disc.
@@ 6.2. Fredholm Theory for Broken Disks
The collection of broken holomorphic disks form a moduli space. We will setup the required Fredholm theory to cut out the moduli space of broken disk as a zero set of a Fredholm section. Since we have punctures going to cylindrical ends, we use weighted Sobolev spaces to create Banach spaces of maps.
Let $\Sigma$ be a punctured disk with punctures at $\mathcal{E}(\Sigma)$. For each puncture $p\in\mathcal{E}(\Sigma)$ we choose a Reeb chord $\gamma_{p}$ or orbit depending on whether $p$ is on boundary or interior. Equip $\Sigma$ with cylindrical holomorphic charts $\phi_{p}$ (i.e. domain of the chart is $[0,1]\times[0,\infty)$ or $S^{1}\times[0,\infty)$ ) near each puncture $p$. We define the set of model maps as follows,
[…]
$u(\phi_{p}(s,t))=(ks,\gamma_{p}(kt))\forall p\in\mathcal{E}(\Sigma)\bigg{\}}$
where $\text{where }k\text{ is the action of }\gamma_{p}$
which consists of all continuous maps which are equal to trivial strip/cylinder in each of the cylindrical end of $\Sigma$.
Select a cylindrical Riemannian metric $g$ on ×$\mathbb{R}\times Z$ and a cylindrical connection $\nabla$ on the tangent bundle. Equip $\Sigma$ with a metric which is of cylindrical form on the charts at each puncture. For every end $p\in\mathcal{E}(\Sigma)$ we will allow Sobolev sections which exponentially decay to parallel transport of sections over the asymptotic Reeb chord or orbit at the puncture, $\eta(p)\in\Omega^{0}(\gamma_{p}^{*}(TZ)\oplus\mathbb{R})$.
Note that for any $u\in\text{Maps}^{0}(\Sigma,\mathbb{R}\times Z,\mathbb{R}\times\Lambda_{2})$ which we write as $(u_{\mathbb{R}},u_{Z})$ using the standard projection maps in ×$\mathbb{R}\times Z$ , there is a natural splitting
@@
where the exponentiation is taken with respect to a metric for which ×$\mathbb{R}\times\Lambda_{2}$ is totally geodesic. We use geodesic exponentiation to define charts of the space $\text{Maps}_{k,p,\delta}$ to give it a structure of Banach manifold.
@@
By a combinatorial type $\Gamma$ of a disk, we mean the data of interior marked points $\mathcal{I}$, puncture node $\mathcal{E}$ , and by ¯$\overline{\mathcal{M}}_{\Gamma}$ we denote the compactified moduli space of disk domains of combinatorial type $\Gamma$. The interior marked points will be the points which get mapped to a Donaldson divisor $D$, see Definition 6.17. We can cover $\mathcal{M}_{\Gamma}$ by charts $\mathcal{M}_{\Gamma_{i}}$ for $i\in\mathcal{A}$ such that the universal moduli space $\mathcal{U}_{\Gamma_{i}}$ above them admit trivializations. For each chart $\mathcal{M}_{\Gamma_{i}}$, we can construct a base space
@@
$ev_{int}:\mathcal{F}^{-1}(0)\longrightarrow M^{|\mathcal{I}|}$
$ev_{int}(u)=\Pi_{m\in\mathcal{I}}\{u(z_{m})\}$
[…]
@@
$ev_{rc}:\mathcal{F}^{-1}(0)\longrightarrow\Lambda_{2}^{|\mathcal{E}|}$
$ev_{rc}(u)=\Pi_{e\in\mathcal{E}}\{\gamma_{e}(0))\}$
@@
In our setup we will be concerned with counting rigid broken disks which pass through a point fixed in $L_{out+}.$ We call a broken disk to be once boundary marked if there is exactly one marking on the boundary of $\mathbb{D}_{out+}.$ Fix $p\in L_{out+}$, a once boundary marked broken disk with a point constraint $p$ would be the moduli space of once marked broken disks which map the boundary marking to $p$. The virtual dimension $\operatorname{vir}_{const}$ of this moduli space is $\operatorname{vir}-\dim(L)$ where the $-\dim(L)$ is the contribution from constraining the marked point to map to $p$.
@@
We will now define the notion of domain dependent perturbation of almost complex structure. $\mathcal{M}_{\overline{\Gamma}}$ to be the moduli space of nodal disk with combinatorial type $\overline{\Gamma}$ and $\mathcal{U}_{\overline{\Gamma}}\xrightarrow{\widetilde{\pi}_{\overline{\Gamma}}}\mathcal{M}_{\overline{\Gamma}}$ is the universal moduli space whose fiber over $S\in\mathcal{M}_{k}$ is the set by pairs $(S,z)$ where $z$ is a point on $S$. Fix an almost complex structure $J_{fix}$ on each piece of a broken manifold obtained by neck stretching. We denote $\mathcal{J}_{cyl}(\widehat{M}_{out\pm},\omega,D)$ to be the set of almost complex structures satisfying the following,
@@
A domain dependent almost complex structure of a disk with $k$ interior markings is a smooth map $\mathcal{J}_{\overline{\Gamma}}:\mathcal{U}_{\overline{\Gamma}}\rightarrow\mathcal{J}_{cyl}(\widehat{M}_{i},\omega,D)$ such that near the puncture nodes, the almost complex structure is equal to $J_{fix}$, i.e., there is an open neighborhood $U$ in $\mathcal{U}_{\overline{\Gamma}}$ near the locus of puncture nodes such that $\mathcal{J}_{\overline{\Gamma}}(w)=J_{fix}$ for $w\in U$
@@
For notational simplicity we illustrate the gluing construction precisely in the case where there is only one puncture node, the general construction is done similarly. Fix a parameter $T>0$. For now, assume $\Gamma^{st}$ has only one puncture node. Let $\Gamma^{st}_{gl}$ be the combinatorial type obtained from gluing $\Gamma^{st}$ across a puncture node. Thus, the moduli space of disks $\mathcal{M}_{\Gamma^{st}}$ of type $\Gamma^{st}$ is a lower dimensional stratum in the compactified moduli space ¯$\overline{}\mathcal{M}_{\Gamma^{st}_{gl}}$. We can choose a chart of $\mathcal{M}_{\Gamma^{st}_{gl}}$ near the strata $\mathcal{M}_{\Gamma^{st}}$ by first choosing strip-like parameterization of the surfaces near the puncture node and then gluing the strips. Let $[\mathfrak{s}]\in\mathcal{M}_{\Gamma^{st}}$, and $C_{[\mathfrak{s}]}$ be the surface corresponding to $[\mathfrak{s}]$, let $p$ be the puncture node lying between the components $C_{-},C_{+}$ of $C_{[\mathfrak{s}]}$. Let $U_{\pm}$ be neighborhoods of $p$ in $C_{\pm}$ such that we can choose charts ϕ±subscriptitalic-$\phi_{\pm}$ on them as follows,
@@
[…]
We explain the notion of being holomorphic with respect to a domain dependent choice of almost complex structure. A map $u:S_{\textbf{z}}\rightarrow\widehat{M}_{out}$ from a $k$ marked disk $S_{\textbf{z}}=\widetilde{\pi}_{k}^{-1}([\textbf{z}])$ where $[\textbf{z}]\in\mathcal{M}_{k}$ is called holomorphic with respect to the domain dependent almost complex structure $\mathcal{J}$ if it satisfies the domain-dependent version of the Cauchy Riemann equation
@@
We can define the following map by setting it to ¯$\overline{\partial}_{j(m),J_{i}}$ on each component
¯$\overline{\partial}_{Br}:\mathcal{M}_{\Gamma^{st},l}\times\text{Maps}^{k,p}_{\delta}(\Sigma_{in},\mathbb{C}^{n},T_{\gamma})\times\text{Maps}^{k,p}_{\delta}(\Sigma_{1},\mathbb{R}\times S^{2n-1},\mathbb{R}\times\Lambda)$
×$\times\dots\text{Maps}^{k,p}_{\delta}(\Sigma_{out},\widehat{M}_{out+},L_{out+},K)\rightarrow$ $\mathcal{E}_{out-}\times\dots\mathcal{E}_{out+}.$
The zero set ¯$\overline{\partial}_{Br}^{-1}(0)$ is the space of holomorphic maps from each piece of the broken disk to the corresponding level of a broken manifold without the matching of Reeb asymptotes. The moduli space of broken disks is obtained as the inverse image of the diagonal of space of Reeb chord matchings,
@@
We say a broken disk $(\mathfrak{u},\mathfrak{D})$ is regular if the linearization of ¯$\overline{\partial}_{Br}$ is surjective at $\mathfrak{u}$ and the evaluation map $EV$ is transverse to $\Delta\times D_{out-}^{\mathcal{I}_{out-}}\times\dots D_{out+}^{\mathcal{I}_{out+}}$ at $EV(\mathfrak{u}).$
@@ 6.4. Regularity of broken buildings
The moduli space of unparametrized interior disk with a single boundary puncture asymptotic to a minimal Reeb chord is of dimension $n-1$ from Lemma 4.7 which is same as the dimension of the space of Reeb chords, $\mathcal{R}(\Lambda_{2})$. We denote this unparameterized moduli space as $\mathcal{M}^{up}_{\Gamma_{in}}(M_{in})$ Thus, choosing a perturbation scheme to cut out the matching condition at each Reeb chord between the inside and outside piece of rigid broken disk ensures that the Reeb chords which occur in rigid broken disks are regular values under the evaluation map from interior pieces.
@@ 6.5. Disk-Potential from SFT buildings
where $u_{g}$ is obtained by a gluing of the broken disk $\mathfrak{u}$. Note that the [$\partial u_{g}$] does not depend on the gluing parameter because of exponential convergence to trivial strips near each puncture.

### v2

@@ 4.1. Background on Neck-Stretching
Definition 4.2 (Double-stretched manifold).
@@
where $\sim^{1}$ is the pair of identifications $\phi(\{a^{\prime}-\epsilon\}\times Z)=\{-T\}\times Z$ and $\phi(\{a^{\prime}+\epsilon\}\times Z)=\{T\}\times Z$ and $\sim^{2}$ is the similar pair of identifications at $\{b^{\prime}\pm\epsilon\}\times Z$. We define a T-double-stretched almost complex structure $J^{T}$ by extending the cylindrical almost complex structure naturally to $T,T]\times Z$.
@@
$\widehat{M}_{out-},\underbrace{M_{-l}:=\mathbb{R}\times Z,\cdots,M_{-1}:=\mathbb{R}\times Z}_{l\text{ times}},{M}_{in}:=\mathbb{R}\times Z,\underbrace{M_{1}:=\mathbb{R}\times Z,\cdots,M_{k}:=\mathbb{R}\times Z}_{k\text{ times}},\widehat{M}_{out+}$
where $\widehat{M}_{out}$ is obtained by attaching natural symplectization ends to $M\setminus\phi([a,b]\times Z).$ The manifolds $\mathbb{R}\times Z$, between $M_{in}$ and $\mathcal{M}_{out\pm}$ are called the neck pieces. If they are above or below the inside piece $M_{in}$, they are referred to as lower or upper neck pieces.
In case $Z$ was a separating hypersurface, we call the two disjoint components of $\widehat{M}_{out}$ to be $\widehat{M}_{out-}$ and $\widehat{M}_{out+}$ depending on which end of symplectization is added.
@@
If $u_{n}:\mathbb{D}^{2}\rightarrow(M^{n},J^{n},L^{n})$ is a sequence of $J^{n}$ holomorphic disc to the $n$-double-stretched manifolds, with boundary on $L^{n}$, and with bounded symplectic energy, i.e. $E(u)<E<\infty$, then there is a subsequence of $u_{n}$ that converges to a stable broken disc.
@@ 4.2. Fredholm Theory for Broken Disks
We will setup the required Fredholm theory to cut out the moduli space of broken disks as a zero set of a Fredholm section. Since we have punctures going to cylindrical ends, we use weighted Sobolev spaces to create Banach spaces of maps.
Let $\Sigma$ be a punctured disk with punctures at $\mathcal{E}(\Sigma)$. For each puncture $p\in\mathcal{E}(\Sigma)$ we choose a Reeb chord $\gamma_{p}$ or orbit depending on whether $p$ is on boundary or interior. Equip $\Sigma$ with cylindrical holomorphic charts $\phi_{p}$ (i.e. domain of the chart is $0,1]\times[0,\infty)$ or $S^{1}\times[0,\infty)$ ) near each puncture $p$. We define the set of model maps as follows,
$\text{Maps}^{0}(\Sigma,\mathbb{R}\times Z,\mathbb{R}\times\Lambda_{2})=\Big\{$ $u\in C^{0}(\Sigma,\mathbb{R}\times Z)\ \Big|\ u(\partial\Sigma)\subset\mathbb{R}\times\Lambda_{2},$
$u(\phi_{p}(s,t))=(ks,\gamma_{p}(kt))\ \text{for all }p\in\mathcal{E}(\Sigma)\,\Big\},$
$\text{where }k\text{ is the action of }\gamma_{p}.$
This set consists of all continuous maps which are equal to trivial strips/cylinders in each of the cylindrical end of $\Sigma$.
Select a cylindrical Riemannian metric $g$ on $\mathbb{R}\times Z$ and a cylindrical connection $\nabla$ on the tangent bundle. Equip $\Sigma$ with a metric which is of cylindrical form on the charts at each puncture. For every end $p\in\mathcal{E}(\Sigma)$ we will allow Sobolev sections which exponentially decay to parallel transport of sections over the asymptotic Reeb chord or orbit at the puncture, $\eta(p)\in\Omega^{0}(\gamma_{p}^{*}(TZ)\oplus\mathbb{R})$.
Note that for any $u\in\text{Maps}^{0}(\Sigma,\mathbb{R}\times Z,\mathbb{R}\times\Lambda_{2})$ which we write as $(u_{\mathbb{R}},u_{Z})$ using the standard projection maps in $\mathbb{R}\times Z$ , there is a natural splitting
@@
where the exponentiation is taken with respect to a metric for which $\mathbb{R}\times\Lambda_{2}$ is totally geodesic. We use geodesic exponentiation to define charts of the space $\text{Maps}_{k,p,\delta}$ to give it a structure of Banach manifold.
@@
By a combinatorial type $\Gamma$ of a disk, we mean the data of interior marked points $\mathcal{I}$, puncture node $\mathcal{E}$ , and by $\overline{\mathcal{M}}_{\Gamma}$ we denote the compactified moduli space of disk domains of combinatorial type $\Gamma$. The interior marked points will be the points which get mapped to a Donaldson divisor $D$, see Definition 4.17. We can cover $\mathcal{M}_{\Gamma}$ by charts $\mathcal{M}_{\Gamma_{i}}$ for $i\in\mathcal{A}$ such that the universal moduli space $\mathcal{U}_{\Gamma_{i}}$ above them admit trivializations. For each chart $\mathcal{M}_{\Gamma_{i}}$, we can construct a base space
@@
$\mathrm{ev}_{\mathrm{int}}:$ $\mathcal{F}^{-1}(0)\longrightarrow M^{|\mathcal{I}|},$
$u\longmapsto\big(u(z_{m})\big)_{m\in\mathcal{I}},$
$\text{where }z_{m}\text{ denotes the }m\text{-th interior marked point.}$
@@
$\mathrm{ev}_{\mathrm{rc}}:$ $\mathcal{F}^{-1}(0)\longrightarrow\Lambda_{2}^{|\mathcal{E}|},$
$u\longmapsto\big(\gamma_{e}(0)\big)_{e\in\mathcal{E}},$
@@
In our setup we will be concerned with counting rigid broken disks which pass through a point fixed in $L_{out+}.$ We call a broken disk to be once boundary marked if there is exactly one marking on the boundary of $\mathbb{D}_{out+}.$ Fix $p\in L_{out+}$, a once boundary marked broken disk with a point constraint $p$ would be the moduli space of once marked broken disks which map the boundary marking to $p$. The virtual dimension $\operatorname{vir}_{const}$ of this moduli space is $\operatorname{vir}-\dim(L)$ where the $\dim(L)$ is the contribution from constraining the marked point to map to $p$.
@@
We will now define the notion of domain dependent perturbation of almost complex structure. Let $\mathcal{M}_{\overline{\Gamma}}$ be the moduli space of nodal disk with combinatorial type $\overline{\Gamma}$. Let $\mathcal{U}_{\overline{\Gamma}}\xrightarrow{\widetilde{\pi}_{\overline{\Gamma}}}\mathcal{M}_{\overline{\Gamma}}$ be the universal moduli space whose fiber over $S\in\mathcal{M}_{k}$ is the set of pairs $(S,z)$ where $z$ is a point on $S$. Fix an almost complex structure $J_{fix}$ on each piece of a broken manifold obtained by neck stretching. We denote $\mathcal{J}_{cyl}(\widehat{M}_{out\pm},\omega,D)$ to be the set of almost complex structures satisfying the following:
@@
A domain dependent almost complex structure of a disk with $k$ interior markings is a smooth map $\mathcal{J}_{\overline{\Gamma}}:\mathcal{U}_{\overline{\Gamma}}\rightarrow\mathcal{J}_{cyl}(\widehat{M}_{i},\omega,D)$ such that near the puncture nodes, the almost complex structure is equal to $J_{fix}$, i.e., there is an open neighborhood $U$ in $\mathcal{U}_{\overline{\Gamma}}$ near the locus of puncture nodes such that $\mathcal{J}_{\overline{\Gamma}}(w)=J_{fix}$ for $w\in U$ .
@@
For notational simplicity we illustrate the gluing construction precisely in the case where there is only one puncture node, the general construction is done similarly. Fix a parameter $T>0$. For now, assume $\Gamma^{st}$ has only one puncture node. Let $\Gamma^{st}_{gl}$ be the combinatorial type obtained from gluing $\Gamma^{st}$ across a puncture node. Thus, the moduli space of disks $\mathcal{M}_{\Gamma^{st}}$ of type $\Gamma^{st}$ is a lower dimensional stratum in the compactified moduli space $\overline{}\mathcal{M}_{\Gamma^{st}_{gl}}$. We can choose a chart of $\mathcal{M}_{\Gamma^{st}_{gl}}$ near the strata $\mathcal{M}_{\Gamma^{st}}$ by first choosing strip-like parameterization of the surfaces near the puncture node and then gluing the strips. Let $\mathfrak{s}]\in\mathcal{M}_{\Gamma^{st}}$, and $C_{[\mathfrak{s}]}$ be the surface corresponding to $\mathfrak{s}]$, let $p$ be the puncture node lying between the components $C_{-},C_{+}$ of $C_{[\mathfrak{s}]}$. Let $U_{\pm}$ be neighborhoods of $p$ in $C_{\pm}$ such that we can choose charts $\phi_{\pm}$ on them as follows,
@@
We call the set
$\mathcal{P}=\{\mathcal{J}_{\overline{\Gamma}}\mid\overline{\Gamma}\text{ is a combinatorial type of broken disk}\}$
, a coherent family of perturbation data if the $T$-glued domain dependent almost complex structure forms a coherent family of perturbation for $M^{T}$ as in Definition 3.14 in [cwstab] for any $T>0$.
We explain the notion of being holomorphic with respect to a domain dependent choice of almost complex structure. A map $u:S_{\textbf{z}}\rightarrow\widehat{M}_{out}$ from a $k$ marked disk $S_{\textbf{z}}=\widetilde{\pi}_{k}^{-1}([\textbf{z}])$ where $\textbf{z}]\in\mathcal{M}_{k}$ is called holomorphic with respect to the domain dependent almost complex structure $\mathcal{J}$ if it satisfies the domain-dependent version of the Cauchy Riemann equation
@@
We can define the following map by setting it to $\overline{\partial}_{j(m),J_{i}}$ on each component
$\overline{\partial}_{Br}:\mathcal{M}_{\Gamma^{st},l}\times\text{Maps}^{k,p}_{\delta}(\Sigma_{in},\mathbb{C}^{n},T_{\gamma})\times\text{Maps}^{k,p}_{\delta}(\Sigma_{1},\mathbb{R}\times S^{2n-1},\mathbb{R}\times\Lambda)$
$\times\dots\text{Maps}^{k,p}_{\delta}(\Sigma_{out},\widehat{M}_{out+},L_{out+},K)\rightarrow$ $\mathcal{E}_{out-}\times\dots\mathcal{E}_{out+}.$
The zero set $\overline{\partial}_{Br}^{-1}(0)$ is the space of holomorphic maps from each piece of the broken disk to the corresponding level of a broken manifold without the matching of Reeb asymptotes. The moduli space of broken disks is obtained as the inverse image of the diagonal of space of Reeb chord matchings,
@@
We say a broken disk $(\mathfrak{u},\mathfrak{D})$ is regular if the linearization of $\overline{\partial}_{Br}$ is surjective at $\mathfrak{u}$ and the evaluation map $EV$ is transverse to $\Delta\times D_{out-}^{\mathcal{I}_{out-}}\times\dots D_{out+}^{\mathcal{I}_{out+}}$ at $EV(\mathfrak{u}).$
@@ 4.4. Regularity of broken buildings
The moduli space of unparametrized interior disk with a single boundary puncture asymptotic to a minimal Reeb chord is of dimension n1n-1 from Lemma 5.7 which is same as the dimension of the space of Reeb chords, $\mathcal{R}(\Lambda_{2})$. We denote this unparameterized moduli space as $\mathcal{M}^{up}_{\Gamma_{in}}(M_{in})$ Thus, choosing a perturbation scheme to cut out the matching condition at each Reeb chord between the inside and outside piece of rigid broken disk ensures that the Reeb chords which occur in rigid broken disks are regular values under the evaluation map from interior pieces.
@@ 4.5. Disk-Potential from SFT buildings
where $u_{g}$ is obtained by a gluing of the broken disk $\mathfrak{u}$. Note that the $\partial u_{g}$] does not depend on the gluing parameter because of exponential convergence to trivial strips near each puncture.